OpenStudy (anonymous):

\[\sec^2 (\pi/7) + \sec^2 (3\pi/7) + \sec^2(5\pi/7)\]

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

I was thinking of making a cubic just for the sake for it.

OpenStudy (anonymous):

\[= \sec^2(2\pi/7) + \sec^2(4\pi / 7) + \sec^2 (6\pi/7)\]Seems better this way.

OpenStudy (anonymous):

Yeah, I think I can make a cubic now...